For the function f(x) = sin x, show with the aid elementary formula


sinA=1/2(1–cos2A)


f(x+y) – f(x) = cos x sin y – 2sin x sin^2 (1/2y)


1
Expert's answer
2021-09-21T11:29:29-0400

Solution:

(1) We know that

"\cos2A=1-2\sin^2A
\\ \Rightarrow 2\sin^2A=1-\cos2A \ ...(i)
\\ \Rightarrow \sin A=\sqrt{\dfrac{{1-\cos 2A}}{2}}"

(2) Given "f(x)=\sin x"

Now, "f(x+y) – f(x)=\sin(x+y)-\sin x"

"=\sin x \cos y+\cos x \sin y-\sin x
\\=\cos x \sin y-\sin x (1-\cos y)
\\=\cos x \sin y-\sin x (2\sin^2\frac y2) \ [using \ (i)]
\\=\cos x \sin y-2\sin x \sin^2\frac y2"